OpenStudy (starryknight):

What is the solution to the following system of equations? -8x + 8y = -16 2x + 6y = -28 please help me( please break it down in steps please)

OpenStudy (***[isuru]***):

lets take -8x + 8y = -16 ------ (1) and 2x + 6y = -28 -----(2) NOW lets multiply (2) by 4 u will get (2) x 4 8x + 24y = - 112----- (3) then add (3) + (1) -8x + 8y + 8x + 24y = -16 + (-112) -8x will cancel out +8x and it will give u 32y = -128 y = -128/32 y = -4 now lets plug (-4) instead of y in (1) -8x + 8y = -16 ------ (1) -8x + 8(-4) = -16 -8x -32 = -16 -8x = 16 x = 16/(-8) x = -2 so.... x= (-2) and y = (-4) hope this will help ya :)

OpenStudy (opcode):

Have you made an attempt? Where are you lost? I will provide an example and steps so you can learn how to do this. The general steps are to solve one of the equations for either "x =" or "y =". So say you have: 3y - 2x = 11 y + 2x = 9 3y - 2x = 11 y = 9 - 2x So replace the "y" value in the first equation by what "y" now equals. 3(9 - 2x) - 2x = 11 Solve this new equation for "x". (27 - 6x) - 2x = 11 27 - 6x - 2x = 11 27 - 8x = 11 -8x = -16 x = 2 Place this new "x" value into either of the ORIGINAL equations in order to solve for "y". y + 2x = 9 or y = 9 - 2x y = 9 - 2(2) y = 9 - 4 y = 5 So we know that: x = 2 and y = 5 You can now check your answer, but substituting in the new found variables. 3y - 2x = 11 3(5) - 2(2) = 11 15 - 4 = 11 11 = 11 \(\huge\color{red}☑\) y + 2x = 9 5 + 2(2) = 9 5 + 4 = 9 9 = 9 \(\huge\color{red}☑\)